Opened 12 years ago

Closed 10 years ago

Last modified 3 years ago

#120 closed flight processing (fixed)

GB07/04, flight day 044/2008, Hayton

Reported by: mark1 Owned by: mark1
Priority: alpha 4 high Milestone: 2008 data processing completion
Component: Processing: general Keywords:
Cc: Other processors:

Description (last modified by mark1)

Data location: ~arsf/arsf_data/in_progress/2008/flight_data/uk/GB07_04-2008_044_Hayton

Data arrived from ARSF via initial SATA disk transfer in March.

Scientific details: See ~arsf/arsf_data/in_progress/2007/ARSF_Applications-GB_2007/GB07-04_*

PI: P. Halcon

Sensors:

  • Eagle (delivered 2nd June 2008)
  • Hawk (delivered 2nd June 2008)

Change History (12)

comment:1 Changed 12 years ago by mggr

  • Summary changed from GB2007/04, flight day 044/2008, Hayton to GB08/04, flight day 044/2008, Hayton

comment:2 Changed 12 years ago by mggr

  • Summary changed from GB08/04, flight day 044/2008, Hayton to GB07/04, flight day 044/2008, Hayton

Oops, right idea, wrong date..

comment:3 Changed 12 years ago by mark1

Broken nav data for Hawk line 3. Processed using -sYS switch, header file and -so to get time offset of 14.65 seconds. This has then been added onto the GPS Start Time in the header file (the original header file has been copied and renamed *hdr_orig).

comment:4 Changed 11 years ago by mark1

  • Description modified (diff)
  • Owner changed from mggr to mark1

comment:5 Changed 11 years ago by mark1

  • Component changed from ARSF to Processing: general
  • Type changed from task to flight processing

comment:6 Changed 11 years ago by mark1

  • Owner changed from mark1 to benj

Delivery directory created and located at
~mark1/scratch_space/GB07_04-2008_044_Hayton/delivery/20080529/

comment:7 Changed 11 years ago by benj

  • Owner changed from benj to mark1

Couple of minor changes, rsynced to ~arsf/arsf_data/in_progress/2008/flight_data/uk/GB07_04-2008_044_Hayton

comment:8 Changed 11 years ago by mark1

  • Description modified (diff)

Data delivered on 2nd June 2008

comment:9 Changed 11 years ago by mark1

PI confirmed, by email, arrival of data on 3rd of June 2008

comment:10 Changed 10 years ago by mggr

  • Resolution set to fixed
  • Status changed from new to closed

Data sent to NEODC for archiving.

comment:11 Changed 6 years ago by tipo

Reprocessed nav and hyperspectral.

comment:12 Changed 3 years ago by dac

Reprocessed hyperspectral data delivered to Rachel Opitz via FTP as not available at NEODC.

Note: See TracTickets for help on using tickets.